OpenStudy (anonymous):

How do i Solve . . . x^2+15x+14

OpenStudy (anonymous):

(x+1)(x+14)

OpenStudy (anonymous):

(x + 14)(x + 1) = 0 x + 14 = 0 x = -14 x + 1 = 0 x = -1 x = -14, -1

OpenStudy (anonymous):

you factor it out, 1 and 14 multiply to 14 and add to 15